the utility function for a consumer utility is U=30Q1 1/2 Q2 1/2. if the price per unit of Q1 =shs 10 and shs 5 per unit of Q2 determine the quantities the consumer should maximize utility if the consumer budgeted shs 350


1
Expert's answer
2022-03-08T09:48:51-0500

The utility function for a consumer utility is "U = 30Q_1^{\frac{1}{2}} Q_2^{\frac{1}{2}}". if the price per unit of Q1 =shs 10 and shs 5 per unit of Q2. Determine the quantities the consumer should maximize utility if the consumer budgeted shs 350.


"U = 30Q_1^{\frac{1}{2}} Q_2^{\frac{1}{2}}"

The budget line is given by:


"350=10Q_1+5Q_2"


When the bundle is optimum, the slope of the budget line matches the slope of the utility function.

"Slope\ of\ BL=\frac{-price\ Of\ Q_2}{price\ of\ Q_1}=\frac{-2}{10}=-0
.5"


"Slope\ of\ U =\frac{-MUQ_1}{MUQ_2}"

"MUQ_1=15Q_2"

"MUQ_2=15Q_1"

"Slope\ of\ U=\frac{-15Q_1}{15Q_2}"


Taking the budget line's slope and applying it to the slope of U:

"\frac{-15Q_1}{15Q_2}=-0.5"

"-15Q_1=-7.5Q_2"

"Q_1=0.52Q_2"


We may achieve the following results by substituting Q1 in the budget line:

"10\u00d7(0.5Q_2)+5Q_2=350"

"10Q_2=350"

"Q_2=35"

"Q_1=0.5Q_2"

"Q_1=17.5"


Therefore, to maximize utility, "Q_1 \u200b= 17.5" and "Q_2=35"